### v1

# Model Extraction Attacks Revisited

Abstract.
Model extraction (ME) attacks represent one major threat to Machine-Learning-as-a-Service (MLaaS) platforms by “stealing” the functionality of confidential machine-learning models through querying black-box APIs. Over seven years have passed since ME attacks were first conceptualized in the seminal work (Tramèr et al., 2016). During this period, substantial advances have been made in both ME attacks and MLaaS platforms, raising the intriguing question: How has the vulnerability of MLaaS platforms to ME attacks been evolving?
In this work, we conduct an in-depth study to answer this critical question. Specifically, we characterize the vulnerability of current, mainstream MLaaS platforms to ME attacks from multiple perspectives including attack strategies, learning techniques, surrogate-model design, and benchmark tasks. Many of our findings challenge previously reported results, suggesting emerging patterns of ME vulnerability. Further, by analyzing the vulnerability of the same MLaaS platforms using historical datasets from the past four years, we retrospectively characterize the evolution of ME vulnerability over time, leading to a set of interesting findings. Finally, we make suggestions about improving the current practice of MLaaS in terms of attack robustness. Our study sheds light on the current state of ME vulnerability in the wild and points to several promising directions for future research.
1. Introduction
The remarkable advances in machine learning (ML) technologies in recent years (He et al., 2016; Deng and Platt, 2014; Roy et al., 2018; Naumov et al., 2019; Feng et al., 2021; Lagler et al., 2013; Zhou et al., 2023) have spurred the expansion of Machine-Learning-as-a-Service (MLaaS), which meets the increasing need for ML capabilities among users who might not possess the requisite expertise or infrastructure. MLaaS platforms offer publicly accessible APIs, enabling users to interface with backend ML models and systems seamlessly. Users are often charged on a per-query basis, making advanced ML capabilities both accessible and affordable. Many IT giants (e.g., Google, Amazon, Microsoft, and Face) have unveiled their MLaaS platforms.
However, there exists an inherent conflict between the proprietary nature of ML models and the public accessibility of MLaaS APIs, leading to a range of security concerns. One of the most prominent concerns is the threat of model extraction (ME) attacks. As illustrated in Figure 1, such attacks construct a piracy model functionally equivalent or similar to the backend model of MLaaS, via querying the black-box APIs with carefully crafted inputs. Since its conceptualization in (Tramèr et al., 2016), a stream of work has been proposed to improve the efficacy of ME attacks (Tramèr et al., 2016; Orekondy et al., 2019a; Krishna et al., 2019; Chandrasekaran et al., 2020; Pal et al., 2019, 2020; Pengcheng et al., 2018; Shi et al., 2018; Juuti et al., 2019; Jagielski et al., 2020; Yu et al., 2020; Truong et al., 2021; Carlini et al., 2021; Papernot et al., 2016a, b, 2018, 2017).
Despite the plethora of prior work, our understanding of the vulnerability of real-world MLaaS to ME attacks remains limited. With the exception of (Yu et al., 2020), most of these studies simulate ME attacks in controlled environments, which may differ significantly from real-world MLaaS scenarios. For instance, some studies (e.g., (Krishna et al., 2019)) assume the adversary possesses prior knowledge about the victim model’s architecture, while others (e.g., (Truong et al., 2021)) create synthetic data to extract models, but MLaaS will initially reject the random noise images, impeding progress. Moreover, since the conceptualization of ME attacks, substantial advances have been made in ME attacks, ML techniques, and MLaaS platforms. It is unclear whether the conclusions drawn in prior work still hold in this rapidly evolving landscape. Specifically, we have the following intriguing questions:
RQ1 - How vulnerable are today’s real-world MLaaS APIs with respect to ME attacks?
RQ2 - How might the adversary exploit such vulnerability effectively (e.g., query-cost reduction)?
RQ3 - How has the vulnerability of MLaaS platforms been evolving in the past years?
Our work – To answer these questions, we design, implement, and evaluate MeBench, the first open-source platform for evaluating the ME vulnerability of MLaaS APIs in a unified and holistic manner. Currently, MeBench has integrated 4 representative ME attacks, 4 attack performance metrics, as well as a suite of 11 piracy models and 6 benchmark datasets. Further, MeBench has implemented a rich set of analysis tools for characterizing the vulnerability, including comparing vulnerability across different APIs, measuring attack transferability, and tracing vulnerability evolution. Our findings can be summarized as follows.
Previous Conclusion |
Refined Conclusion |
Consistency |
Identifying the architectures of victim models and using the same architectures in piracy models substantially boost the effectiveness of ME attacks (Chen et al., 2022b). |
The architectures of piracy models have a limited impact on the performance of ME attacks, while more advanced models may not lead to more effective attacks. |
Using more complex models tends to achieve better attack performance (Shi et al., 2017; Krishna et al., 2019; Aïvodji et al., 2020). |
The impact of model complexity varies with the concrete tasks (i.e., FER versus NLU) while pre-training is a more dominating factor. |
Semi-supervised learning substantially improves the query efficiency of ME attacks, especially when the query budget is low (Jagielski et al., 2020). |
Semi-supervised learning improves the query efficiency of ME attacks, but the margin of improvement is not as large as reported in local experiments (Jagielski et al., 2020); further, it has a negative impact on adversarial fidelity. |
Active learning substantially improves the query efficiency of ME attacks (Pal et al., 2020). |
Active learning yields only marginal improvements in query efficiency and can, in some cases, have a negative effect. |
Using adversarial examples improves both the query efficiency and the attack effectiveness (Papernot et al., 2017; Juuti et al., 2019; Pengcheng et al., 2018; Yu et al., 2020). |
Using adversarial examples has a limited or even negative impact on attack fidelity, but may improve adversarial fidelity. |
Perturbing output confidence scores effectively mitigates ME attacks (Orekondy et al., 2019b; Lee et al., 2019). |
Output quantization weakens ME attacks but is not sufficiently effective. |
– Leveraging MeBench, we conduct an empirical study on leading MLaaS platforms (i.e., Amazon, Microsoft, Face, and Google) in the tasks of facial emotion recognition (FER) and natural language understanding (NLU). We show that today’s real-world MLaaS APIs still exhibit significant vulnerability to ME attacks, while the characteristics of vulnerability vary greatly across different platforms and tasks.
– We further examine the influential factors on the performance of ME attacks, including optimizers, training regimes, model architectures, and advanced attack strategies. Our evaluation leads to a set of interesting findings, many of which challenge the conclusions in prior work, as summarized in Table 1. For instance, it is found that compared with other factors (e.g., optimizer), the piracy models have a limited impact on the attack performance, challenging the conventional notion that more advanced models lead to more effective attacks. Also, it is shown that adversarial examples offer little boost to or may even negatively impact the efficacy of ME attacks, which contradicts the findings in prior work.
– Finally, by integrating a longitudinal dataset containing over 1.7 million queries to leading MLaaS platforms spanning from 2020 to 2022, we conduct a retrospective study to characterize the vulnerability evolution of these MLaaS platforms. We discover some significant trends in vulnerabilities and identify the impact of model updates on ME attacks. Our analysis leads to some notable findings, including the influence of model updates on ME attack results and the potential lack of investment in model protections. Our contribution highlights the historical perspective of the evolving ME attack in MLaaS platforms, emphasizing the necessity for enhanced security measures and proactive defense against ME attacks.
We envision that the MeBench platform and our findings facilitate future research on ME attacks and shed light on building MLaaS platforms in a more secure manner.
2. Background
2.1. Preliminaries
We first introduce fundamental concepts and assumptions used throughout the paper.
Deep neural networks (DNNs) (LeCun et al., 2015) represent a class of ML models to learn high-level abstractions of complex data. In a predictive task, a DNN (parameterized by ) encodes a function , which maps an input to a class . The training of often involves iteratively updating via algorithms such as stochastic gradient descent (SGD) (Robbins and Monro, 1951), aiming to minimize a loss function (e.g., cross-entropy) that measures the discrepancy between the model’s prediction and the ground-truth class . Techniques such as data augmentation, batch normalization, and dropout may also be used during training to accelerate training or prevent overfitting. As their design may require significant engineering effects and their training may involve substantial data and compute resources, DNN models are often considered invaluable intellectual property in various contexts.
Knowledge distillation (KD) (Hinton et al., 2015; Gou et al., 2021) is a process where a student model is trained to mimic a teacher model . Typically, KD involves minimizing the discrepancy between the two models, which can be measured in responses (i.e., models’ outputs) (Hinton et al., 2015), features (i.e., models’ intermediate representations) (Huang and Wang, 2017; Romero et al., 2014), or relations (i.e., models’ modeling of input relationships) (Lee et al., 2018; Yim et al., 2017). Formally,
where measures the discrepancy between two models with respect to a reference dataset .
Model extraction (ME) aims to infer a victim model’s properties typically through black-box query access. The inferred information may include the model’s architecture, (hyper)parameters, functionality, and other properties (e.g., attack vulnerability). The existing ME attacks can be roughly categorized as exact or approximate extraction.
Exact extraction aims to infer the victim model’s properties exactly, for instance, architectures (Oh et al., 2019), hyperparameters (Wang and Gong, 2019), and parameters (Tramèr et al., 2016) (with respect to known architectures). For instance, equation-solving attacks (Tramèr et al., 2016) allow the extraction of the exact parameters of (multi-class) logistic regression and multi-layer perceptron models. Approximate extraction aims to construct a piracy model similar to the victim model (Yu et al., 2020; Pal et al., 2020; Tramèr et al., 2016; Papernot et al., 2016b; Gong et al., 2021), which can be further divided based on its goal: 1) obtaining similar performance as the victim model (measured by accuracy) or 2) acquiring similar behavior as the victim model (measured by fidelity).
In the following study, we primarily focus on KD-based, approximate ME attacks, due to their general applicability and limited assumptions.
2.2. Threat model
We first define the threat model assumed in our study.
Adversary’s objective – The adversary’s goal is to generate a piracy model that approximates the behavior and/or performance of the MLaaS backend model . The agreement between and is measured by either accuracy or fidelity with respect to a testing dataset .
Adversary’s knowledge – We assume a black-box setting, in which the adversary has little knowledge about the victim model , including its architecture, (hyper)parameters, and the specific dataset used in its training. Yet, the adversary has access to a reference (unlabeled) dataset .
Adversary’s capability – The adversary is able to access through the MLaaS API, which, for a given query input , returns ’s prediction . Note that the queries are not restricted to real data and may include synthesized or adversarial inputs. Moreover, as MLaaS often charges users on a per-query basis, we assume the adversary has a limited query budget .
2.3. A general attack framework
In MeBench, all the ME attacks are implemented within a general framework, as illustrated in Figure 2. This framework comprises three key functions i) GenerateQuery, ii) QueryAPI, and iii) UpdateModel. Specifically, GenerateQuery generates queries that are either sampled from the reference dataset or synthesized; QueryAPI then sends each query to the MLaaS API and receives its prediction ; finally, UpdateModel optimizes the piracy model using the query-response pairs . Notably, this process can be executed iteratively: GenerateQuery generates queries based on both the updated piracy model and the previous query results (i.e., using active learning to identify the next batch of queries).
In the default ME attack, we sample inputs from to query the MLaaS API and use all the query-response pairs to train the piracy model by minimizing the following cross-entropy loss:
3. Status quo of ME vulnerability
Leveraging MeBench, we conduct an empirical study on leading MLaaS platforms. Our study is designed to center around 4 questions:
Q: How vulnerable are today’s MLaaS APIs to ME attacks?
Q: How do various factors influence such vulnerability?
Q: How does the vulnerability evolve over time?
Q: How effective are the defenses (if any) employed on the MLaaS platforms?
3.1. Experiment setting
We begin by introducing the setting of our evaluation.
Datasets – To be succinct yet reveal key issues, we primarily use 6 benchmark datasets: RAFDB (Li and Deng, 2019), EXPW (Zhang et al., 2016), KDEF (Lundqvist et al., 2022), FER+ (Barsoum et al., 2016), IMDB (imd, [n. d.]), and YELP (yel, [n. d.]), spanning both the vision and natural language domains. The first 4 correspond to the facial emotion recognition (FER) task, which classifies given facial images into 7 possible expressions (e.g., “calm”). The last 2 correspond to the natural language understanding (NLU) task, which classifies given text into 4 sentiments (i.e., “positive”, “negative”, “neutral”, and “mixed”). The details of the datasets are deferred to Table 18. By fault, following prior work, we partition each dataset into an 80%/20% split, designating the 80% as the reference data (for ME attacks) and the remaining 20% as the testing data (for performance evaluation). In addition, we also evaluate the scenario in which and come from different datasets in §3.4.
APIs – We consider the APIs of 4 leading MLaaS service providers: Amazon, Google, Microsoft, and Face. By default, we assume the adversary has access to the complete query response (i.e., classification labels and confidence scores). Specifically, in the FER task, each query response contains the confidence scores of different expressions (e.g., “calm”). In the NLU task, each query response includes the confidence scores of different sentiments (e.g., “positive”). The details of each API are deferred to Table 14.
Piracy models – In FER, we consider 4 representative architectures for the piracy model: VGG (Simonyan and Zisserman, 2014), ResNet (He et al., 2015), DenseNet (Huang et al., 2017), and ViT (Dosovitskiy et al., 2020), while in NLU, we consider 2 Transformer-based architectures for the piracy model: RoBERTa (Liu et al., 2019) and XLNet (Yang et al., 2019). Using models of distinct architectures (e.g., residual blocks versus skip connects), we aim to factor out the influence of individual model characteristics. Besides the backbone model, we use one fully connected layer with softmax activation as the classification head. By default, we assume the piracy models are randomly initialized and trained from scratch. In §3.3.2, we also consider the scenario in which the piracy models are pre-trained on public datasets. In the FER task, the models are pre-trained on the ImageNet-1K dataset; in the NLU task, the models are pre-trained on the BERT dataset (Devlin et al., 2018).
Metrics – To evaluate the effectiveness of ME attacks, we mainly use three metrics:
Accuracy measures the fraction of inputs in the testing set that are correctly classified by the piracy model . Formally,
where denotes the indicator function that returns 1 if the predicate is true and 0 otherwise.
Fidelity measures the fraction of inputs in the testing set that receive the same classification by the victim and piracy models (Yu et al., 2020; Jagielski et al., 2020; Pal et al., 2019; Correia-Silva et al., 2018; Papernot et al., 2017). Formally,
Adversarial fidelity measures the fraction of adversarial examples (with respect to ) that receive the same classification by and . Formally,
where is the adversarial attack (e.g., PGD (Madry et al., 2018)) that generates adversarial examples with respect to . We mainly use this metric in adversarial ME attacks (§3.3.3).
| API | Dataset | |||
| KDEF | RAFDB | EXPW | FER+ | |
| Amazon | +0.490.03 | -0.980.15 | -1.230.092 | +1.990.12 |
| Microsoft | / | +1.230.19 | +2.120.17 | +4.120.18 |
| Face | +2.120.18 | +1.780.13 | +1.970.04 | +0.450.14 |
Data augmentation – In the vision domain, data augmentation (e.g., random cropping and resizing) plays a significant role in model training. Thus, we conduct a pilot study on its effectiveness in the basic ME attack, with results shown in Table 2. Observe that data augmentation improves the attack fidelity in most cases. Thus, we apply data augmentation by default in the FER experiments.
ME attacks – We evaluate a variety of ME attacks. The basic attack is based on knowledge distillation (Yim et al., 2017), which randomly samples inputs from the reference dataset as queries, receives query response from the MLaaS API for each query , and subsequently trains the piracy model based on the query-response pairs . Moreover, we take into account more advanced attacks that aim to minimize the number of queries through various strategies (e.g., semi-supervised learning) in §3.3.3.
3.2. Q1: Overall vulnerability
3.2.1. Measurements
We first examine the overall vulnerability of MLaaS APIs to ME attacks, with results summarized in Figure 3. We have the following observations.
Platform variations – A detailed evaluation of different APIs on the RAFDB dataset during ME attacks shows significant variations. The attack fidelity varies from 63.53% against the Amazon API to 83.98% against the Microsoft API. These variations can be attributed to several factors:
(i) Label discrepancies – For instance, Amazon employs unique 8-class labels, diverging from the typical 7-class labels used by other platforms and the dataset itself.
(ii) Training data distributions – While we lack precise knowledge about the original training data across different platforms, this factor is highly likely to influence the outcomes of ME attacks.
(iii) Pre-processing and post-processing – The unique pre-processing (e.g., data cleaning, normalization, and augmentation) and post-processing (e.g., confidence score quantization) implemented by different MLaaS platforms may also have a substantial impact on the ME vulnerability.
Task/dataset variations – Our findings also show that the ME vulnerability varies across tasks and datasets. This is evident in the significant fluctuation in the attack fidelity across different tasks and datasets. We attribute these variations to the unique characteristics intrinsic to each dataset.
In the NLU task, take the Amazon API as an example. Recall that the output of Amazon API falls into 4 categories (i.e., “positive”, “negative”, “mixed”, and “neutral”), where “mixed” exists as a separate class. IMDB is a high-polarity dataset, while Yelp is a diverse semantic dataset including a number of instances with mixed sentiments. This unique feature of the Yelp dataset allows the attack to better capture the “mixed” class, leading to higher attack fidelity compared with the IMDB dataset.
Compared with the NLU task, the FER task requires classifying given facial expressions into 7/8 emotion classes, without ambiguous “mixed” classes. This explains the relatively lower attack fidelity in the FER task. Further, observe that against the Amazon API, the attack fidelity on KDEF is much higher than the other datasets. This is explained by that KDEF explicitly instructs subjects to exhibit exaggerated expressions during data collection, which makes the classification task much easier. However, the piracy model extracted from KDEF does not generalize to other datasets (more details in §3.4).
Fidelity vs. accuracy – Figure 3 also shows that the variations in fidelity (63% to 85%) and accuracy (49% to 77%) across different APIs do not always align with each other. For instance, in the FER task, the attack against the Amazon API attains the highest fidelity yet the lowest accuracy on KDEF across different datasets. This is explained by that the agreement between the victim and piracy models, measured by fidelity, is not necessarily correlated with their performance, measured by accuracy. Given that ME attacks aim to extract the functionality of victim models, rather than attaining high performance, similar to prior work (Gong et al., 2021; Orekondy et al., 2019a; Pal et al., 2020; Yu et al., 2020), we primarily focus on the metric of attack fidelity in the study below.
The findings above highlight the need for effective mitigation in mainstream MLaaS APIs. Moreover, it reiterates the importance of comprehending different factors (e.g., datasets and tasks) for accurately interpreting the ME vulnerability across different MLaaS platforms.
3.2.2. Comparison with prior work
We also compare our results with prior work on ME vulnerability in both settings of local models and MLaaS APIs.
Local models – Most prior work (e.g., (Gong et al., 2021; Orekondy et al., 2019a; Pal et al., 2020)) focuses on ME attacks against local models and uses CIFAR10 as the primary dataset. In particular, InverseNet (Gong et al., 2021) shows superior performance over the other ME attacks on CIFAR10. If the adversary is able to access the full confidence information, InverseNet attains 45%, 70%, 81%, and 82% attack fidelity under 1K, 5K, 10K, and 15K queries, respectively. The fidelity growth plateaus after 15K queries, indicating a diminishing return from further increasing the query budget.
In Figure 3, the Microsoft API shows the highest vulnerability, with 84% attack fidelity for EXPW (27K queries) and RAFDB (12K queries), respectively. Due to its smaller size, the lower number of queries for FER+ (3K queries) results in slightly lower fidelity than the other two datasets. These findings corroborate with prior work, showing a positive positive correlation between query budget and attack fidelity.
MLaaS APIs – Among prior work, Cloudleak (Yu et al., 2020) conducts ME attacks against MLaaS APIs and measures attack fidelity with respect to different APIs and query budgets. Specifically, it reports 81% fidelity and 58% accuracy against the Face API on the KDEF dataset. Furthermore, it reports that the attacks against other platforms such as Microsoft and Clarifai, as well as different tasks including traffic sign recognition, flower, and NSFW, all attain over 82% fidelity and 70% accuracy. This clearly showcases the widespread nature of this vulnerability across various APIs and tasks.
In our experiments, on the same platform and dataset (i.e., KDEF and Face), we observe that the attack attains similar results (74% fidelity and 66% accuracy), while the performance gap with (Yu et al., 2020) may be attributed to the scratch model and the difference of testing data. In (Yu et al., 2020), the testing data is synthesized, while we randomly sample 20% of the KDEF dataset as the testing set. Further, similar to (Yu et al., 2020), the ME vulnerability varies considerably across different APIs. In particular, the attack attains over 80% fidelity against the Microsoft API across all the datasets, while the other APIs demonstrate more variations from one dataset to another.
Overall, we may conclude:
3.3. Q2: Influential factors
Next, we evaluate how different key factors impact the effectiveness of ME attacks with respect to given MLaaS APIs.
3.3.1. Optimizers
Prior studies indicate that the selection of optimizers greatly impacts training dynamics (Wu et al., 2021). Thus, we examine how it affects ME attack performance.
We specifically consider 4 popular optimizers: SGD (Robbins and Monro, 1951), Adam (Kingma and Ba, 2017), AdamW (Loshchilov and Hutter, 2019), and Lion (EvoLved Sign Momentum) (Chen et al., 2023), which is a recently proposed optimizer. We configure the basic ME attack with varied optimizers and evaluate its effectiveness in the FER and NLU tasks.
| API | Model | Optimizer | |||
| Lion | AdamW | Adam | SGD | ||
| RAFDB in FER | |||||
| Amazon | ResNet50 | 62.890.53 | 64.150.17 | 63.960.06 | 56.060.65 |
| Microsoft | 84.640.20 | 84.350.11 | 84.350.10 | 81.010.65 | |
| IMDB in NLU | |||||
| Amazon | XLNet | 84.300.16 | 82.450.20 | 82.670.32 | 71.690.27 |
| RoBERTa | 81.200.21 | 74.720.08 | 74.580.12 | 42.930.32 | |
| Microsoft | XLNet | 86.350.23 | 86.570.52 | 86.570.24 | 84.490.36 |
| RoBERTa | 86.550.51 | 85.200.54 | 85.150.36 | 65.750.29 | |
FER – We compare different optimizers using RAFDB on the Amazon and Microsoft APIs. Following (Chen et al., 2023), we set the hyper-parameters (e.g., learning rate and decay rate ) for different optimizers as follows: SGD with and , Adam/AdamW with and , and Lion with and . We have the following key observations.
Table 3 shows the attack fidelity at the end of 200 training epochs. Under the same query budget, Lion, Adam, and AdamW attain comparable attack fidelity. There is no distinct advantage between Lion and AdamW; however, it is evident that SGD lags behind the performance of other optimizers significantly. We further examine the training dynamics of different optimizers. Figure 4 summarizes how the attack fidelity varies with the number of training epochs. Despite the claimed superior convergence speed of Lion in supervised learning (Chen et al., 2023), we observe that Adam/AdamW actually converges much faster than Lion in the ME attack, while SGD converges fairly slowly.
NLU – We also set the hyper-parameters of optimizers following (Chen et al., 2023): SGD with and , Adam/AdamW with and , and Lion with and .
Table 3 summarizes the attack performance corresponding to different optimizers on the IMDB dataset. Lion generally outperforms the other optimizers on both Microsoft and Amazon APIs. Our observations corroborate prior findings (Chen et al., 2023) that Lion outperforms alternative optimizers in fine-tuning large language models.
Overall, we may conclude:
3.3.2. Piracy models
The architectures of piracy models are also crucial for ME attacks. It is shown in prior work (Chen et al., 2022b) that correctly identifying the architectures of victim models and using the same architectures in piracy models substantially boost the effectiveness of ME attacks. Several studies (Orekondy et al., 2019a; Shi et al., 2017; Krishna et al., 2019; Aïvodji et al., 2020) also demonstrate that the piracy model needs to be at least as complex as the victim model. Additionally, the results of (Shi et al., 2017; Krishna et al., 2019; Aïvodji et al., 2020) suggest that utilizing a more complex model leads to better attack performance.
| Component | Model | Pre-training | #Params (M) | Model Family |
| Model Family | GoogLeNet | / | 6.7 | GoogLeNet |
| DenseNet | 8 | DenseNet | ||
| EfficientNet | 56 | EfficientNet | ||
| AlexNet | 61 | AlexNet | ||
| ResNet50 | 25.6 | ResNet | ||
| Model Complexity | ResNet18 | / | 11.7 | ResNet |
| ResNet50 | 25.6 | |||
| ResNet101 | 44.6 | |||
| WResNet50 | 68.9 | |||
| Use of Pre-training | ResNet50 | / | 25.6 | ResNet |
| ImageNet1K | ||||
| VGGFace2 |
In the open-world setting targeted in this study, the victim models are not accessible, limiting us from directly comparing the architectures of piracy and victim models. Therefore, in the study below, we evaluate how different aspects of piracy models may impact the ME attacks with respect to given MLaaS APIs. Specifically, we explore three key aspects of model architectures: model families, model complexity (within the same family), and (non-)use of pre-training. Table 4 summarizes the experimental settings in evaluating the impact of piracy models.
Model family – For FER, we compare 5 popular architectures, GoogLeNet (Szegedy et al., 2014), DenseNet (Huang et al., 2018), ResNet (He et al., 2016), EfficientNet (Tan and Le, 2020), Alex -Net (Krizhevsky, 2014), using RAFDB on the Microsoft and Face APIs, with results shown in Figure 5. For NLU, we compare 2 popular LLMs, RoBERTa and XLNet, using IMDB on the Amazon and Microsoft APIs, with results summarized in Table 5. We have the following interesting findings.
| Model | Variant | API | |
| Amazon | Microsoft | ||
| XLNet | Base | 84.50 | 86.19 |
| Large | 84.30 | 86.35 | |
| RoBERTa | Base | 78.96 | 86.12 |
| Large | 81.20 | 86.62 | |
While prior work suggests that the piracy model architecture greatly impacts the effectiveness of ME attacks (e.g., 10-30% difference in attack fidelity (Chen et al., 2022b)). However, in our evaluation, except for AlexNet, an architecture proposed in 2012, which substantially under-performs the other architectures, the difference among other architectures is relatively marginal. A similar phenomenon is also observed in the NLU task as shown in Table 5.
It is worth pointing out that the number of parameters in these architectures varies from 6.7K to 61K. Yet, interestingly, GoogleLeNet, with only 6.7K parameters, delivers performance comparable with more complex architectures. To further evaluate the impact of model families, we consider ViT (Dosovitskiy et al., 2020), a more advanced architecture with 86K parameters. However, it proves challenging to optimize ViT due to the limited data; even with pre-training, ViT does not match ResNet in terms of attack fidelity. This implies that for ME attacks, overly complex architectures can be counterproductive.
Model complexity – We further assess the impact of model complexity (within a single model family). In FER, we use ResNet as the piracy model and vary its width and depth, spanning across ResNet18, ResNet50, ResNet101, and Wide-ResNet50. As illustrated in Figure 6, the attack fidelity across different architectures differs by less than and on Amazon and Face, respectively. Intuitively, as the model complexity reaches a certain level, further increasing the model width or depth does not lead to a notable improvement in attack performance. Similarly, in NLU, as shown in Table 5, for both RoBERTa and XLNet, using large models yields marginal improvement over base models. It is worth pointing out that compared with the base models, the large models of RoBERTa and XLNet are pre-trained using more data ( additional data) (Yang et al., 2019; Liu et al., 2019).
Use of pre-training – Next, we evaluate the influence of pre-training piracy models using public datasets on the performance of ME attacks.
| Model | Pre-training | API | |
| Amazon | Microsoft | ||
| XLNet-Base | ✗ | 36.39 | 65.78 |
| ✓ | 84.50 | 86.19 | |
| RoBERTa-Base | ✗ | 53.94 | 73.61 |
| ✓ | 78.96 | 86.11 | |
In FER, we compare the attacks using a randomly initialized ResNet50 model as the piracy model with that using ResNet50 models pre-trained on public datasets. We consider two pre-training datasets, ImageNet-1K and VggFace2 (Cao et al., 2018). As shown in Figure 7, pre-training marginally improves the attack performance, while the improvement margin varies with the APIs. For instance, compared with VGGFace2, ImageNet-1K leads to more improvement on Microsoft but less so on Face. In NLU, we compare randomly initialized XLNet and RoBERTa and that pre-trained on the BERT dataset (Devlin et al., 2018), with results shown in Table 6. Surprisingly, unlike the PER task, pre-training language models significantly improves the attack performance. For instance, the pre-training of RoBERTa boosts fidelity by over 25%.
Thus, we may conclude:
| API | Dataset | Attack | Type | Query Budget (# Batches) | |||||||
| 4 | 8 | 16 | 32 | 64 | 128 | Full | |||||
| Microsoft | RAFDB | Basic | Fid | 68.940.37 | 73.80.88 | 77.360.13 | 79.430.33 | 81.420.27 | 83.290.16 | 84.350.13 (191 Batches) | +1.12 |
| MixMatch | 70.251.35 | 75.010.56 | 78.670.16 | 80.620.18 | 82.470.35 | 83.930.02 | |||||
| EXPW | Basic | 70.251.15 | 73.420.42 | 76.981.7 | 78.450.72 | 80.170.19 | 81.540.31 | 83.940.08 (417 Batches) | +1.44 | ||
| MixMatch | 72.41(1.03) | 74.881.2 | 77.651.33 | 80.050.57 | 81.60.03 | 82.850.41 | |||||
| Amazon | FER+ | Basic | Fid | 41.410.56 | 48.091.22 | 53.840.20 | 59.660.32 | / | / | 63.200.41 18.571.12 (50 Batches) | +5.91 |
| AdvFid | 18.160.45 | 16.490.22 | 17.261.60 | 18.590.61 | / | / | |||||
| MixMatch | Fid | 49.731.42 | 55.391.40 | 59.21.15 | 62.320.16 | / | / | ||||
| AdvFid | 16.620.74 | 16.051.20 | 16.070.67 | 15.71.07 | / | / | |||||
3.3.3. Attack strategies
Thus far, we mainly focus on the basic attack strategy that randomly samples queries from the reference dataset. Next, we investigate advanced strategies that aim to minimize the number of queries. Specifically, we consider three popular strategies: semi-supervised learning, active learning, and adversarial learning.
Semi-supervised learning – Via semi-supervised learning, the adversary queries the API with a small number of samples and mixes the query responses (labeled data) and unlabeled data to train the piracy model. We consider MixMatch (Berthelot et al., 2019) as a representative semi-supervised learning method, which, in the context of ME attacks, is employed to reduce the number of queries (Jagielski et al., 2020). At a high level, MixMatch-based ME attack samples a batch of samples to query the API as the labeled data and mixes it with an equally-sized batch of unlabeled data to produce a batch of augmented labeled data and a batch of augmented unlabeled data with pseudo labels , which are used to train the piracy model but with different objective functions.
We evaluate the MixMatch-based ME attack in the FER task. Specifically, using the RAFDB and EXPW datasets on the Microsoft API, we compare the performance of basic and MixMatch-based ME attacks with varying numbers of queries, ranging from 4 to 128 batches (with batch size fixed as 64). Note that the FER+, RAFDB, and EXPW datasets include 50, 191, and 417 batches, respectively. As summarized in Table 7, we have the following interesting findings.
(i) MixMatch generally enhances the effectiveness of ME attacks across different datasets and APIs. However, the improvement is less significant than that reported in (Jagielski et al., 2020). For instance, it improves the fidelity by about 1% against the Microsoft API and about 5% against the Amazon API, under varying query budgets. The divergence from prior work may be attributed to both models and data. For instance, the results in (Jagielski et al., 2020) are evaluated using a one-layer fully-connected network on the CIFAR10 data, while the MLaaS backend models often use much more complex architectures trained on a massive amount of data (e.g., millions of images) (Amazon, [n. d.]).
(ii) Intriguingly, while boosting attack fidelity, MixMatch has a negative impact on adversarial fidelity, which measures the agreement between the victim and piracy models with respect to adversarial examples. For instance, the basic attack achieves adversarial fidelity 2.9% higher than the MixMatch-based attack with batches. This observation may be explained as follows. MixMatch augments the training data by interpolating different samples, which essentially “smooths” the decision boundaries. While improving query efficiency, it also prevents the piracy model from effectively learning the high-curvature decision boundaries of the victim model, which is essential for adversarial fidelity.
Active learning – The adversary may also employ active learning to choose informative inputs rather than randomly sampled ones to query the API to reduce the query cost.
We adopt -Center-Greedy (Sener and Savarese, 2017), a method also used in Active -Thief (Pal et al., 2020), in ME attacks. At each iteration, ActiveThief picks central samples from the unlabeled data (determined as cluster centers in the piracy model ’s latent space) to query the API. These responses are added to the training set , which is used to train . This process continues until the allocated query budget is exhausted. The results are summarized in Table 8. We have the following findings.
| API | Dataset | Attack | Type | Query Budget (# Batches) | |||||
| 16 | 32 | 64 | 96 | 128 | Full | ||||
| Microsoft | RAFDB | Basic | Fid | 76.740.76 | 79.280.54 | 81.760.08 | 82.740.15 | 83.20.39 | 84.350.13 (191 Batches) |
| ActiveThief | 76.450.31 | 79.170.4 | 82.020.33 | 82.930.3 | 840.16 | ||||
| EXPW | Basic | 75.940.56 | 78.520.45 | 79.760.08 | 81.080.11 | 81.570.24 | 83.940.08 (417 Batches) | ||
| ActiveThief | 74.560.68 | 77.090.42 | 79.420.09 | 80.430.47 | 81.040.21 | ||||
| Amazon | RAFDB | Basic | Fid | 53.520.57 | 57.070.61 | 59.410.72 | 61.080.8 | 61.690.41 | 63.530.14 22.810.13 (191 Batches) |
| AdvFid | 21.610.01 | 21.290.02 | 22.521.06 | 22.20.95 | 23.320.41 | ||||
| ActiveThief | Fid | 52.140.07 | 56.030.28 | 58.810.36 | 60.320.7 | 61.811.17 | |||
| AdvFid | 20.340.6 | 21.260.98 | 22.050.8 | 22.990.95 | 22.550.31 | ||||
In our experiment combination of Microsoft+RAFDB, the results from Active Learning appear inconsistent. Some seed and query budgets show slight improvements, but these results are unstable, especially with lower query budgets. For instance, Active Learning performs worse with query budgets of 16 and 32 batches than the Basic attack. With a budget of 64 batches, there is a slight average improvement of 0.26%. However, as the query budget rises, the performance becomes steadier, although the improvement is never over 1%. This initial inconsistency might be attributed to the initial model struggles to capture face feature maps well. When the model is capable, the number of queries has been exhausted. As a result, when the model becomes more capable, the number of available queries has already been exhausted, which may prevent it from selecting the best data points. We notice a similar pattern with the Amazon+RAFDB combination. AdvFid and Fid perform worse than the Basic attack at low query budgets. Otherwise, in the Microsoft+EXPW combination, Active Learning always performs worse than the Basic attack.
Back to previous research on active learning, when -Center-Greedy combines with data augmentation (Mittal et al., 2019), results in only slight accuracy improvements. Specifically, they show an increase of approximately 1 to 2% on CIFAR10 compared to random sampling-based queries. In earlier studies related to ME attacks, (Pal et al., 2020) demonstrates that using the -Center-Greedy method leads to an average improvement of 0.976% in attack fidelity on CIFAR10. However, when the query budget increases to 25K, this method under-performs by -0.62%. In previous work, it is effective in most cases but has limited improvement and shows instability.
In conclusion, active learning-based attacks on different APIs and datasets often perform worse than basic ME attacks. They demonstrate lower fidelity and adversarial fidelity, especially when there are fewer data queries. Given its inconsistent performance, active learning is not recommended for ME attacks in real-world scenarios.
Adversarial learning – CloudLeak (Yu et al., 2020) applies adversarial learning to generate examples to reduce the required number of queries in ME attacks. Specifically, the strategy is to perform adversarial attacks on clean data and produce adversarial examples on the piracy model . These examples represent highly uncertain regions of . When these regions are queried against the API, they offer valuable insights to refine . Additionally, since lies close to ’s decision boundaries, these queries push towards the decision boundaries of the victim model . We consider two representative adversarial attacks: PGD (Madry et al., 2018) (with , , , and random initialization) and CW (with , , and ) (Carlini and Wagner, 2017). The results are summarized in Table 9.
| API | Dataset | Query Batches | Attack | ||
| Basic | PGD | CW | |||
| Amazon | RAFDB | 8 | 40.460.39 | 36.970.1 | 37.50.22 |
| 17.270.27 | 19.020.05 | 20.350.35 | |||
| 16 | 43.480.47 | 41.660.17 | 40.390.45 | ||
| 19.660.17 | 19.160.29 | 21.270.15 | |||
| 32 | 45.943.81 | 42.922.02 | 46.912.94 | ||
| 21.330.15 | 20.710.22 | 23.40.48 | |||
| FER+ | 8 | 35.321.83 | 39.910.2 | 41.521.58 | |
| 17.060.01 | 18.860.97 | 18.311.28 | |||
| 16 | 41.51.83 | 42.330.02 | 46.43.97 | ||
| 19.020 | 20.810.08 | 17.754.91 | |||
| 32 | 44.560.2 | 49.350.38 | 50.290.32 | ||
| 17.870.74 | 19.461.34 | 17.280.35 | |||
| Face | RAFDB | 8 | 40.920.41 | 40.921.27 | 35.072.7 |
| 20.310.29 | 20.790.09 | 20.560.14 | |||
| 16 | 44.653.25 | 40.010.19 | 41.090.07 | ||
| 18.580.34 | 16.570.48 | 17.150.05 | |||
| 32 | 47.910.46 | 45.280.18 | 45.480.29 | ||
| 15.540.21 | 16.20.43 | 17.290.24 | |||
| FER+ | 8 | 67.790.31 | 22.581.8 | 51.020.01 | |
| 14.460.2 | 23.980.46 | 18.740.3 | |||
| 16 | 72.613.39 | 45.60.49 | 41.652.02 | ||
| 14.460.46 | 18.280.2 | 13.20.2 | |||
| 32 | 67.790.13 | 47.510.01 | 47.510.27 | ||
| 14.460.47 | 17.790.01 | 17.570.22 | |||
It is observed that against the Amazon API, using PGD or CW to enhance ME attacks leads to a considerable improvement in attack fidelity (around 5.61% on average) on FER+, while the effect on RAFDB is fairly mixed. Meanwhile, against Face, adversarial learning-enhanced ME attack seems not only ineffective but even under-performs the basic attack. However, across all the cases, adversarial learning leads to a consistent improvement in adversarial fidelity. Our findings indicate that clean queries (sampled from the reference dataset) and adversarial queries (generated by adversarial attacks) seem to contribute to ME attacks differently, respectively improving the fidelity and adversarial fidelity of piracy models.
| Dataset | Type | Adversarial/Clean Ratio | ||||
| 1:0 | 3:1 | 1:1 | 1:3 | 0:1 | ||
| RAFDB | Fid | 42.92 | 44.85 | 43.78 | 45.41 | 45.94 |
| AdvFid | 20.71 | 20.57 | 20.55 | 20.37 | 21.33 | |
| FER+ | Fid | 49.35 | 44.2 | 43.14 | 47.00 | 44.71 |
| AdvFid | 19.46 | 18.62 | 20.06 | 20.07 | 17.35 | |
These observations lead us to explore an interesting question: whether combining clean and adversarial inputs in querying the MLaaS API improve the overall efficacy of ME attacks. To investigate this, we vary the ratio of clean and adversarial (generated by PGD) queries in each batch while keeping the query budget constant. The experimental results are summarized in Table 10.
In some cases (e.g., RAFDB), including adversarial queries, may negatively impact the attack fidelity. In other cases (e.g., FER+), although only using adversarial queries leads to the best fidelity, the proportion of adversarial queries is not consistently correlated with the attack fidelity, which conflicts with the findings in (Yu et al., 2020). Similarly, although using more adversarial queries improves adversarial fidelity, increasing the proportion of adversarial queries is not strictly correlated with the increase of adversarial fidelity.
It is possible to explain these observations as follows. There exists intricate dynamics between clean and adversarial queries: while adversarial queries often represent corner cases with respect to the model’s decision boundaries, clean queries represent more average cases. Thus, while increasing the number of queries generally improves the agreement between the victim and piracy models, clean and adversarial queries have different focuses.
The results highlight the complexity of operating this attack strategy as well as its dependency on concrete datasets and APIs. Based on the findings, further research is needed to comprehend the possible advantages of applying adversarial learning in ME attacks and the optimal strategy to combine clean and adversarial queries to achieve high attack efficacy.
Overall, we may conclude:
3.4. Q3: Generalizability
Generalizability is a crucial metric for ME attacks: it measures how the extracted piracy model agrees with the victim model on datasets other than that used in the attack. In other words, generalizability indicates how valuable and flexible the extracted piracy model is. Next, under the MLaaS setting, we empirically measure the generalizability of piracy models. We extract the piracy model by querying the API using the reference dataset and evaluate ’s fidelity and accuracy on another dataset .
3.4.1. FER
We conduct experiments across four datasets on three APIs to analyze the generalizability of ME attacks in the FER task, with results summarized in Table 11. The highlighted cells indicate when the original and evaluation datasets are identical, serving as a baseline. The results under Microsoft+KDEF are unavailable due to the closure of the Microsoft API. We have the following key observations from Table 11.
| Evaluation Dataset | API | Origin Dataset | |||
| KDEF | RAFDB | EXPW | FER+ | ||
| KDEF | Amazon | 77.78 | 59.90 | 67.36 | 55.21 |
| Microsoft | / | / | / | / | |
| Face | 73.61 | 63.72 | 67.01 | 50.00 | |
| RAFDB | Amazon | 33.64 | 63.53 | 57.38 | 48.34 |
| Microsoft | / | 85.01 | 83.98 | 75.17 | |
| Face | 37.67 | 66.76 | 66.02 | 57.15 | |
| EXPW | Amazon | 37.35 | 48.03 | 71.09 | 53.14 |
| Microsoft | / | 72.75 | 84.31 | 77.09 | |
| Face | 47.79 | 56.91 | 68.40 | 58.91 | |
| FER+ | Amazon | 33.42 | 49.68 | 65.85 | 63.2 |
| Microsoft | / | 72.80 | 83.13 | 80.36 | |
| Face | 40.78 | 56.06 | 66.39 | 64.29 | |
Asymmetric generalizability – A piracy model extracted from dataset may perform well on dataset , but the reverse is not always true. For instance, the model extracted from the EXPW dataset demonstrates strong generalizability across datasets, especially matching or surpassing baselines on RAFDB and FER+. However, the models extracted from RAFDB and FER+ do not perform as well when applied to EXPW. We speculate that highly generalizable datasets, such as EXPW, often benefit from their larger query volumes, allowing them to capture the distributions of smaller, less adaptable datasets like RAFDB and FER+. Specifically, with its 27K queries, EXPW substantially overshadows RAFDB’s 12K and FER+’s 3K. Given their shared tasks, EXPW’s comprehensive nature likely contributes to its superior generalizability.
Generalizability beyond query budgets – Though KDEF and FER+ boast similar query volumes, KDEF’s generalizability lags significantly. This can be attributed to its unique data collection methodology, where subjects were instructed to exhibit exaggerated emotions. This induces a distribution shift in KDEF, making it deviate from the normative distributions of facial image datasets. Conversely, RAFDB, FER+, and EXPW, derived from normal photos, have distributions that are more congruent with one another.
3.4.2. NLU
Table 12 shows the generalizability results in the NLU task. ME attacks using IMDB as the reference dataset demonstrate strong generalizability, particularly on Microsoft, where fidelity reaches 90.49% compared to the baseline of 92.18% when transferred to Yelp. However, the generalizability is weaker on Amazon, with a fidelity of 80.71% compared to the baseline of 90.88%. The results of Yelp are interesting, showing poor generalizability to Amazon (similar to IMDB) at 68.05%, but high generalizability to Microsoft like IMDB.
In general, the generalizability of ME attacks in NLU is related to the APIs, Amazon is weaker than Microsoft, possibly because Amazon uses 4 classes compared to Microsoft uses 3 classes. Further, IMDB’s data is highly polarized, and Yelp’s data also contains neutral and mixed sentiments, which might also contribute to the results of generalizability.
| Evaluation Dataset | API | Origin Dataset | |
| IMDB | Yelp | ||
| IMDB | Amazon | 84.30 | 68.05 |
| Microsoft | 86.35 | 84.60 | |
| Yelp | Amazon | 80.71 | 90.88 |
| Microsoft | 90.49 | 92.18 | |
3.5. Q4: Defenses
It is unclear what specific defense mechanisms are employed by the MLaaS platforms. However, when examining their APIs, Amazon, Microsoft, and Face will provide exact confidence values for each class. In contrast, Google’s FER categorizes its output using five distinct likelihood descriptors, (i.e., “very_unlikely”, “unlikely”, “possible”, “likely” and “very_likely”) offering a different approach to presenting its analysis. This approach seems to be an attempt to obscure the confidence levels of the model’s output. The central concept behind such quantization is to reduce the granularity of the confidence values returned by the API, thereby making it more challenging for potential attackers to discern intricate details about the model’s internal processes. We mimicked Google’s approach and assess the effectiveness of quantization by applying it to other APIs. Specifically, We achieve this by discretizing the confidence values into intervals and substituting the confidence value with the midpoint of the corresponding interval (e.g., is represented as ). We manually reduce the granularity of the returned confidence information to examine its impact on the performance of the Model Extraction (ME) attack.
| API | Dataset | |||
| KDEF | RAFDB | EXPW | FER+ | |
| Amazon | -1.540.1 | -2.760.11 | -1.980.14 | -0.450.03 |
| Face | +0.430.08 | -0.760.18 | +0.310.12 | -0.980.08 |
Table 13 presents the analysis results. When applied to the Amazon API, quantization resulted in an average decrease of 1.68% in the attack fidelity. In contrast, the Face API exhibits more mixed results in mitigating ME attacks.
Overall, the manual reduction of granularity in the returned confidence information has a somewhat detrimental effect on the results of the ME attack. However, it is not a sufficiently effective defense strategy against model extraction.
4. A retrospective study
To comprehend the evolution of ME vulnerability over the years, we integrate a longitudinal dataset into MeBench and conduct a retrospective study on the ME vulnerability of leading MLaaS APIs, spanning the period from 2019 until now. The purpose of this study is to understand the evolving security risks associated with ME attacks.
4.1. Study setting
In this retrospective study, we employ HAPI (Chen et al., 2022a), a longitudinal compilation containing 1,761,417 queries submitted to commercial MLaaS APIs including Amazon, Google, IBM and Microsoft. The dataset spans from 2020 to 2022 and covers a range of tasks, such as image tagging, speech recognition, and text mining. Each data point comprises a query input along with the MLaaS’s response (i.e., prediction, annotation, and corresponding confidence scores).
Among the range of tasks covered by the HAPI dataset, we focus on two representative tasks due to the inherent constraints of ME: FER (representing computer vision tasks) and NLU (representing natural language processing tasks). Note that the responses of all the APIs include confidence information for all the classes, with the notable exception of Google’s API (details in Table 14).
| Task | API | HAPI1 | Current2 | # Classes |
| FER | Amazon | ✗ | ✓ | 8 |
| Microsoft3 | ✓ | ✗3 | 7 | |
| Face++ | ✓ | ✓ | 7 | |
| ✓ | ✓ | 7 | ||
| NLU | Amazon | ✓ | ✓ | 4 |
| Microsoft | ✗ | ✓ | 3 |
1
“HAPI” represents the data collected by HAPI during 2020-2022 with only the highest confidence scores.
2
“Current” represents the data collected by us in 2023 with full confidence scores across all classes, except for Google API which only returns the highest confidence scores.
3
Microsoft has sunset its FER API for emotion prediction to mitigate potential misuse that subject people to stereotyping, discrimination, or unfair denial of services.
Due to the unique structures of the HAPI dataset, we face a particular challenge: it provides only the highest confidence score among all the classes, without any information about the remaining classes. To address this challenge, we impute the confidence scores of the remaining classes. Specifically, let the prediction about an input be a probability simplex over classes, with each element corresponding to one distinct class. Recall that the API only outputs the highest confidence score, , where . Following the maximum-entropy principle (Jaynes, 1957), which suggests choosing a distribution that retains the most significant degree of uncertainty (or entropy) while adhering to known constraints, we thus assign the confidence scores as:
That is, the class with the highest confidence score is assigned the score given by HAPI, while the remaining classes receive an equal assignment . Note that the sum of confidence scores across all classes equals to 1, ensuring a valid probability distribution.
Also note that in the NLU task, as HAPI only records the highest confidence score between “positive” and “negative”, we expand the outputs to three classes: “positive”, “negative”, and “mixed” in the following study.
4.2. Results
Our study uncovers interesting trends in the evolving vulnerability of MLaaS APIs to ME attacks. Next, we present our findings in both FER and NLU tasks.
4.2.1. FER
Figure 8 measures the attack fidelity across various APIs and datasets from the year 2020 to 2022. Compared with other platforms, there is a general upward trend for the attack fidelity against the Microsoft API across all three datasets, with a significant shift from 2020 to 2021 on the FER+ dataset. To understand the reason behind this shift, we further conduct a detailed analysis, summarizing the variations in the outcomes of ME attacks and the output data of original APIs in Table 15.
| Dataset | ME Attack | API Ouput | ||
| Fidelity | Accuracy | Predicted Class | Avg. Confidence | |
| FER+ | +5.93% | +3.03% | 96.07% | 0.0169 |
| RAFDB | +1.10% | -0.01% | 99.76% | 0.0022 |
| EXPW | +0.38% | +0.03% | 99.44% | 0.0059 |
When examining the API outputs with respect to the HAPI dataset, we find that the predictions in 2020 and 2021 overlap more than 99% for RAFDB and EXPW, while this overlap is only around 96% for FER+. It is important to note that these overlaps only match predicted labels and do not account for variations in confidence scores. When considering the average change in confidence scores, FER+ also exhibits much higher variance than other datasets. Finally, the overall accuracy of the Microsoft API in classifying FER+ also increases by 3.03% from 2020 to 2021, while the accuracy for the other two datasets remains little changed.
The substantial changes in the Microsoft API’s accuracy and confidence scores with respect to the FER+ and EXPW datasets suggest that the backend model may have undergone significant changes from 2020 to 2021. These alterations likely contribute to the varying fidelity of ME attacks. Specifically, the attack attains 5.93% fidelity increase on FER+, with only minor increments of 1.10% and 0.38% on RAFDB and EXPW, respectively.
While ME attacks appear to become both easier and simpler across different datasets, we believe such changes are mainly due to adjustments in the backend model of the Microsoft API (e.g. incorporating new training data to fine-tune the model), rather than any alterations to its defenses against ME attacks. What is particularly interesting is that the significant improvements in model accuracy coincide with the increased attack fidelity. We speculate that after the update, the backend model becomes more aligned with the distribution of the FER+ dataset. As the ME attack specifically targets this dataset, it is easier for the attack to adapt to the backend model. In comparison, Google+EXPW shows an overlap of 90.83% between 2020 and 2021 in HAPI, with a slight fidelity change of -0.47%. All the other dataset-API combinations experience minimal changes, with overlaps consistently exceeding 99%.
In summary, it is evident that in the FER task, the ME vulnerability of various MLaaS APIs has evolved in the past few years. However, we do not observe clear patterns in such evolution and are therefore unable to conclude whether these APIs have strengthened defenses against ME attacks. We tend to believe that most MLaaS platforms have not implemented specific defenses, especially in the case of the Microsoft API, where it exhibits high vulnerability to ME attacks.
4.2.2. NLU
Due to the limitation of NLU datasets in the HAPI dataset, only the combination of Amazon and IMDB is available for analyzing the evolution of ME vulnerability.
| Year | ME Attack | API Ouput | ||
| Fidelity | Accuracy | Predicted Class | Avg. Confidence | |
| 2020-2021 | -2.43% | -1.08% | 83.20% | 0.1682 |
| 2021-2022 | -0.20% | +0.12% | 99.41% | 0.0070 |
As shown in Table 16, the attack fidelity against the Amazon API shows a declining trend over the three years, with a decrease of 2.43% from 2020 to 2021. However, since this change is not statistically significant, it does not appear that the API has fortified its defenses against ME attacks in the intervening years.
Despite the growing popularity of MLaaS APIs, our study indicates a potential lack of investment in safeguarding backend models. This poses questions about the commitment of commercial MLaaS providers to addressing ME threats and implementing protective measures. Considering the ongoing expansion and increasing reliance on MLaaS, it becomes imperative for providers to prioritize security. We advocate for a proactive stance by MLaaS providers to bolster their APIs and fortify the protection of their backend models.
5. Additional related work
Since the concept of ME attacks against MLaaS APIs was first introduced in (Tramèr et al., 2016), which applies path-finding attacks to extract classification (e.g., decision trees) and regression (e.g., regression trees) models, over the years, a plethora of ME attacks have been proposed, training the piracy model to mimic the behavior of the victim model by leveraging data labeled by the victim model (Tramèr et al., 2016; Orekondy et al., 2019a; Krishna et al., 2019; Chandrasekaran et al., 2020; Pal et al., 2019, 2020; Pengcheng et al., 2018; Shi et al., 2018; Juuti et al., 2019; Jagielski et al., 2020; Yu et al., 2020; Truong et al., 2021; Carlini et al., 2021; Papernot et al., 2016a, b, 2018, 2017).
Two key metrics, accuracy and fidelity, have been proposed in (Jagielski et al., 2020) to measure the performance of ME attacks. Fidelity is generally acknowledged as a more critical metric than accuracy for assessing ME attacks, which is also the main metric used in this study. In addition, we also introduce adversarial fidelity, another important metric to assess the agreement between the victim and piracy models with respect to adversarial examples, complementing fidelity and accuracy.
The number of queries is another important metric for ME attacks. Semi-supervised learning has been explored in (Jagielski et al., 2020) to reduce the number of queries. Active learning strategies have also been explored in (Tramèr et al., 2016; Chandrasekaran et al., 2020; Pal et al., 2019, 2020; Pengcheng et al., 2018; Shi et al., 2018). For instance, in (Pal et al., 2020), strategies such as uncertainty, -center, and DeepFool-based active learning are employed to identify the most informative samples. While the experiments span both image and language-based tasks, they are conducted in controlled laboratory environments and not validated on MLaaS platforms. Other work (Papernot et al., 2017; Juuti et al., 2019; Pengcheng et al., 2018; Yu et al., 2020) leverages adversarial examples to optimize the number of queries or improve attack performance. In addition to these studies, which have focused primarily on classification tasks, (Hu and Pang, 2021) explores model extraction attacks on GANs. This work examines whether the conclusions in controlled local environments hold for ME attacks against real-world MLaaS APIs, leading to a number of interesting findings that complement existing studies and provide new insights about the vulnerability of real-world MLaaS APIs.
6. Conclusion
In this paper, we report a systematic study on the vulnerability of real-world machine-learning-as-a-service (MLaaS) APIs to model extraction (ME) attacks. The evaluation leads to a number of interesting findings that complement the existing studies conducted in controlled laboratory environments: (i) Despite their striding progression over the years, leading MLaaS platforms continue to be highly susceptible to ME attacks. (ii) The advances in ML techniques (e.g., optimizers) significantly enhance the adversary’s capabilities. (iii) Attack techniques (e.g., adversarial learning), proven to be effective in controlled environments, may not necessarily exhibit the same level of effectiveness against real-world MLaaS APIs. (iv) The existing mitigation (e.g., output quantization) may weaken ME attacks but remain inadequately effective. These findings shed light on developing MLaaS platforms in a more secure manner.
Appendix A More Details
A.1. Notations
| Notation | Definition |
| , , | generic, online, HAPI victim model |
| piracy model | |
| clean labeled data | |
| clean unlabeled data | |
| adversarial data |
A.2. Datasets
| Task |
Dataset |
Size |
Description |
| F E R |
KDEF |
562*762 |
The exaggerated facial expressions made by subjects |
RAFDB |
100*100 |
Great-diverse facial images downloaded from the Internet |
EXPW |
224*224 |
Wild human facial expressions captured at social events |
FER+ |
48*48 |
Black and white human face expressions |
| N L U |
IMDB |
234 words |
Binary sentiment classification dataset, highly polar movie reviews collected from the internet movie database (IMDB) |
Yelp |
140 words |
Information from user reviews of different restaurants on Yelp. Ratings can be from 1-5, not a polarized dataset |
A.3. Local experiments
To validate the correctness of our experimental setting, we perform local experiments using the CIFAR10 dataset. We train a victim model using the ground-truth labels and use the same strategy in§3.3.3 to perform ME attack against it. The results are summarized in Table 19, 20, and 21.
| Attack | Query Budget (# Batches) | |||||
| 4 | 8 | 16 | 32 | 64 | Full | |
| Basic | 35.05 | 50.41 | 49.46 | 71.98 | 73.92 | 92.88 |
| MixMatch | 49.35 | 58.68 | 79.21 | 79.22 | 83.23 | |
| Attack | Query Budget (# Batches) | |||||
| 8 | 16 | 32 | 64 | 128 | Full | |
| Basic | 47.85 | 48.12 | 68.84 | 76.63 | 83.12 | 92.88 |
| ActiveThief | 52.01 | 58.57 | 74.22 | 80.53 | 86.06 | |
| Query Batches | Attack | ||
| Basic | PGD | CW | |
| 8 | 33.94 | 36.87 | 35.67 |
| 16 | 41.5 | 45.16 | 46.76 |
| 32 | 66.78 | 69.81 | 71.2 |